Which rectangular equation corresponds to these parametric equations?
x = 0.2sec t and y = -0.25tan t
a)25x^2 − 16y^2 = 1
b)25x^2 + 16y^2 = 1
c)x^2 − 16y^2 = 25
d)25y^2 − 16x^2 = 1
e)4x^2 − 16x^2 = 1
I don't even know the first step to this problem- I tried looking up the steps to solving equations like it but i couldnt find any.
Any help is much appreciated! :)

Answer by stanbon(75887)

Which rectangular equation corresponds to these parametric equations?
x = 0.2sec t and y = -0.25tan t
Just substitute the given x and y values into each equation
to see if you get a true statement::
Example::
a)25x^2 − 16y^2 = 1
25(0.2sec(t))^2 - 16(-0.25tan(t))^2 = 1
25(0.04)sec^2(t) - 16(0.0625)tan^2(t) = 1
sec^2(t) - tan^2(t) = 1
(1+tan^2(t)) - tan^2(t) = 1
1 = 1
----------------------------------
Do the same for the following problems.
Cheers,
Stan H.
--------------------
b)25x^2 + 16y^2 = 1
c)x^2 − 16y^2 = 25
d)25y^2 − 16x^2 = 1
e)4x^2 − 16x^2 = 1